# Identification and estimation of vaccine effectiveness in the test-negative design under equi-confounding

### v3

@@
(May 12, 2025)
The test-negative design (TND) is frequently used to evaluate vaccine effectiveness in real-world settings. In a TND study, individuals with similar symptoms who seek care are tested for the disease of interest, and vaccine effectiveness is estimated by comparing the vaccination history of test-positive cases and test-negative controls. Traditional approaches justify the TND by assuming either (a) receiving a test is a perfect proxy for unmeasured health-seeking behavior or (b) vaccination is unconfounded given measured covariates —both of which may be unrealistic in practice. In this paper, we return to the original motivation for the TND and propose an alternative justification based on the assumption of odds ratio equi-confounding, where unmeasured confounders influence test-positive and test-negative individuals equivalently on the odds ratio scale. We discuss the implications of this assumption for TND design and provide alternative estimators for the marginal risk ratio among the vaccinated under equi-confounding, including estimators based on outcome modeling and inverse probability weighting as well as a semiparametric estimator that is doubly-robust. When the equi-confounding assumption does not hold, we suggest a sensitivity analysis that parameterizes the magnitude of the deviation on the odds ratio scale. We conduct a simulation study to evaluate the empirical performance of our proposed estimators under a wide range of scenarios.
@@ 1 Introduction
Post-market evaluation of vaccine effectiveness is critical for understanding how vaccines perform in real-world settings, especially as population immunity changes or pathogens undergo antigenic mutation or selection [patel_postlicensure_2020, hitchings_effectiveness_2021, israel_elapsed_2021]. These studies can also assess effectiveness in subpopulations that were ineligible or underrepresented [olson_effectiveness_2022], evaluate outcomes and safety signals that may have been underpowered [thompson_effectiveness_2021], or compare vaccine formulations not considered in the original trials [skowronski_two-dose_2022]. Often, these evaluations must rely on observational data [chua_use_2020-1, dean_covid-19_2021].
@@
The test-negative design (TND) is a common and cost-efficient study design for estimating vaccine effectiveness [sullivan_potential_2014, jackson_test-negative_2013]. In its canonical form, patients with acute respiratory illness are prospectively screened and recruited based on a unified symptom set through hospitals or other care facilities and tested for the pathogen of interest. Vaccination history and other detailed covariate information are collected at enrollment or through linked medical or vaccination records. Effectiveness is then estimated by comparing the odds of vaccination among test-positive (cases) and test-negative (controls) individuals.
@@
The TND has been previously promoted as a way to eliminate bias due to unmeasured healthcare-seeking behavior — a common source of bias in vaccine evaluations [jackson_test-negative_2013]. Because participants seek a test conditional on the same symptom set but without knowledge of the causative agent, the design effectively restricts to “care-seekers”, thereby precluding it as potential confounder [jackson_test-negative_2013]. Subsequent work pointed out that the TND (i) may not be robust to unmeasured healthcare-seeking when it is nonbinary or when there are other factors affecting vaccine uptake and risk of infection [sullivan_theoretical_2016, lewnard_theoretical_2021, lipsitch_observational_2016] and (ii) is subject to possible selection bias due to conditioning on receiving a test [sullivan_theoretical_2016, lipsitch_observational_2016].
[…]
@@
In this article, we revisit the original motivation and show that the TND may be justified under a less restrictive assumption that unmeasured confounding is equivalent for test-positive cases and test-negative controls on the odds ratio scale. This assumption builds on the insight that infection with another respiratory illness, if unaffected by focal vaccination, is essentially a negative outcome control [lipsitch_negative_2010, shi_selective_2020]. Therefore, any observed difference in the incidence of test negative illnesses between vaccinated and unvaccinated individuals likely reflects residual confounding. Under the assumption of confounding equivalence, we can use the difference in test-negative illness to de-bias our estimate of vaccine effectiveness, similar to the bias correction via the parallel trends assumption in the difference-in-differences literature [sofer_negative_2016, park_universal_2023, tchetgen_universal_2023]. As we show, this approach is valid under the outcome-dependent sampling of the TND, where only those who are tested are sampled. When our assumptions hold, we show that the traditional odds ratio estimator identifies the conditional causal risk ratio among the vaccinated. We also derive new estimators of the marginal risk ratio among the vaccinated, including those based on outcome modeling and inverse probability weighting. We discuss the implications of our assumptions for the design of TNDs and provide proofs of our results as well as simulations to illustrate the consequences when our assumptions are violated. We also propose a sensitivity analysis to evaluate the impact of deviations from equi-confounding on our effect estimates.
@@ 2 Setup and observed data
    the test is perfect
These assumptions can be relaxed (e.g., we discuss imperfect testing in Appendix H). Under this testing regime, we have that $Y=I\times T$ and we consider $Y\neq 0$ to define medically-attended illness, i.e., symptomatic infection severe enough to warrant medical attention. Finally, let $Y^{v}$ denote the potential outcome under $V=v$ for $v=0,1$.
@@ 3 Causal estimands for the test-negative design
In a TND study, the primary estimand is the causal risk ratio,
@@
comparing the probability of medically-attended illness due to the pathogen of interest if everyone were vaccinated versus unvaccinated in the population underlying the TND, with causal vaccine effectiveness (VE) defined as $1-RR$. Occasionally, the causal odds ratio is presented as the target parameter instead of the risk ratio. However, this is often combined with an assumption —either explicit or implicit— that symptomatic infection is rare, in which case the odds ratio approximates the risk ratio.
@@
may nevertheless be point identified. The parameter $\Psi$ is analogous to the average treatment effect on the treated (ATT) in the causal inference literature, though expressed as a ratio rather than a difference. Similarly, we can define a modified VE as $1-\Psi$, interpreted as the proportion of medically-attended illness prevented by vaccination among the vaccinated subjects. When protection is homogeneous, $\Psi$ coincides with the causal risk ratio, i.e. $RR=\Psi$; otherwise, when protection is heterogeneous, $RR$ is a weighted average of the causal risk ratios among the vaccinated and unvaccinated.
@@ 4 Illustrative example
To build intuition, we start with a brief account of the identification of the conditional risk ratio among the vaccinated, $\Psi(X)$, before turning to formal identification results for $\Psi$ in section 5.
@@ 4.1 Decomposition of the conditional causal risk ratio among the vaccinated
This decomposition expresses $\Psi(X)$ as the product of two terms: the (crude) observed risk ratio and a de-biasing term. The de-biasing term quantifies the degree of unmeasured confounding between the vaccinated and unvaccinated on the ratio scale, i.e., the net influence of all unmeasured causes of vaccination and the outcome. When there is no unmeasured confounding, the de-biasing term equals one and the observed risk ratio equals the causal risk ratio among the vaccinated. Conversely, in the presence of unmeasured confounding, the de-biasing term is the quantity necessary to remove bias from the observed risk ratio. However, without additional assumptions, this term (and in particular the denominator) cannot be identified from the observed data.
@@ 4.3 Identification under cohort and outcome-dependent sampling
Nevertheless, under the invariance of the odds ratio [breslow_regression_1976], we can show that $\Psi_{om}(X)$ is equivalent to
@@
which is identified under the sampling design of the TND. As discussed further in section 6, $\Psi^{*}_{om}(X)$ is identical to the conditional odds ratio estimand commonly targeted in TND studies [jackson_test-negative_2013], suggesting that estimates from existing studies could be reinterpreted and justified under the equi-confounding conditions established here.
@@ 5.1 Identifiability conditions
    where $\text{where }OR_{y}(X)=\frac{\Pr[Y^{0}=y|V=1,X]\Pr[Y^{0}=0|V=0,X]}{\Pr[Y^{0}=0|V=1,X]\Pr[Y^{0}=y|V=0,X]}.$
@@
Assumption A1 is a well-known identifiability condition discussed in more detail elsewhere [hernan_causal_2020]. Assumption A2 has two implications. First, it implies that the vaccine does not offer any cross-protection against other types of infection which may cause the same symptoms, including any short-term, nonspecific protection via activation of the immune system. Second, it requires that an individual’s decision to seek care and get tested is unaffected by their vaccination status, though it may depend on other measured and unmeasured factors. In theory, Assumption A2 could be evaluated in a randomized trial by comparing the incidence of testing negative (or testing positive for specific non-target pathogens) across arms. However, in practice, trials may not be adequately powered to rule out small effects. Both aspects of Assumption A2 have been invoked in previous literature on analysis of TND studies [jackson_test-negative_2013, feng_assessment_2017, schnitzer_estimands_2022]. In Appendix F, we show that Assumption A2 can be further relaxed if the effect of vaccination on testing is equivalent for test-positive and test-negative illnesses.
[…]
@@
Finally, Assumption A4 is similar to, although less restrictive than, the well-known positivity condition in causal inference [hernan_causal_2020]. Specifically, Assumption A4 requires overlap of vaccination propensity between the observed vaccinated and unvaccinated among the test positive cases as well as overlap of vaccination propensity across test-positive cases and test-negative controls within observed vaccination groups.
@@ 5.2 Identification in a cohort study
Proof of Proposition 1 is deferred to Appendix B.1. ∎
@@
Following standard terminology [robins_estimating_1992], we refer to Expression 6 as the outcome-modeling estimand and Expression 7 as the inverse-probability weighting estimand.
@@
The quantity $\Pr[V=1|Y^{0}=2,X]$ is sometimes referred to as the extended propensity score (EPS) function, as it extends the standard propensity score to allow for unmeasured confounding by conditioning on the treatment-free potential outcome — in this case $Y^{0}=2$. As shown previously [tchetgen_single_2023], in the presence of unmeasured confounding, knowledge of the EPS is sufficient to identify treatment effects among the treated. Similarly knowledge of $\Pr[V=1|Y^{1}=2,X]$ would permit identification of treatment effects among the untreated, and knowledge of both would permit the identification of the overall effect [tchetgen_single_2023].
@@
[…]
@@ 5.3 Identification under outcome-dependent sampling in a TND
Proof of Proposition 2 is deferred to Appendix B.2. ∎
@@
Expressions 8 and 9 resemble those proposed assuming unconfoundedness [jiang_tnddr_2023]. In that work, the authors introduce a class of marginal estimands in a TND that can be written as
@@ 6 Estimation
We now describe estimation strategies for the causal risk ratio among the vaccinated in a TND, based on the identifying expressions in Proposition 2. However, we describe in Appendix C estimators based on the results in Proposition 1 in a cohort study. For convenience, we define an indicator for test positive case in the TND sample $Y^{*}=\mathbbm{1}(Y=2)$.
@@
where $\pi_{y}(X)=\Pr[V=1\mid S=1,Y^{*}=y,X]$ is the probability of vaccination among the test-negative controls. This probability could be obtained, for instance, via a logistic regression of $V$ on $X$ among those with $Y^{*}=0$. In Appendix C, we discuss methods for estimating standard errors and confidence intervals for $\widehat{\Psi}_{om}^{*}$ and $\widehat{\Psi}_{ipw}^{*}$ via bootstrapping or stacked estimating equations.
The first estimator, $\widehat{\Psi}_{om}^{*}$, requires correct specification of the outcome model, while the second estimator, $\widehat{\Psi}_{ipw}^{*}$, requires correct specification of the extended propensity score model. In practice, however, both models may be difficult to specify correctly. To address this, we construct a doubly-robust estimator for $\Psi$ that is consistent and asymptotically normal if the model for either $\mu^{*}_{0}(X)$ or $\pi^{*}_{0}(X)$ or both are correctly specified. Full details of the doubly robust approach are included in Appendix E, but are based on noting that an equivalent expression of $\Psi_{om}$ and $\Psi_{ipw}$ is
@@
[…]

### v4

@@
(August 11, 2025)
The test-negative design (TND) is frequently used to evaluate vaccine effectiveness in real-world settings. In a TND study, individuals with similar symptoms who seek care are tested for the disease of interest, and vaccine effectiveness is estimated by comparing the vaccination history of test-positive cases and test-negative controls. The design has previously been justified on the grounds that it reduces confounding due to unmeasured health-seeking behavior, although this has not been formally described using potential outcomes. However, it is also widely acknowledged that, by conditioning participation on receipt of a test, the TND risks inducing selection bias. In this paper, we propose a formal justification for the TND based on the assumption of odds ratio equi-confounding, where unmeasured confounders influence test-positive and test-negative individuals equivalently on the odds ratio scale, with health-seeking behavior being just one plausible example. We also show that these results hold under outcome-dependent sampling design of the TND. We discuss the implications of the equi-confounding assumption for TND design and provide alternative estimators for the marginal risk ratio among the vaccinated under equi-confounding, including estimators based on outcome modeling and inverse probability weighting as well as a semiparametric estimator that is doubly-robust. When the equi-confounding assumption does not hold, we suggest a straightforward sensitivity analysis that parameterizes the magnitude of the deviation on the odds ratio scale. We conduct a simulation study to evaluate the empirical performance of our proposed estimators under a wide range of scenarios. Finally, we also discuss how test-negative outcomes may be used more broadly to de-bias estimates from cohort studies where testing is symptom-triggered.
@@ 1 Introduction
Post-market evaluation of vaccine effectiveness is critical for understanding how vaccines perform in real-world settings, especially as population immunity changes or pathogens undergo antigenic mutation or selection [1, 2, 3]. These studies can also assess effectiveness in subpopulations that were ineligible or underrepresented [4], evaluate outcomes and safety signals that may have been underpowered [5], or compare vaccine formulations not considered in the original trials [6]. Often, these evaluations must rely on observational data [7, 8].
@@
The test-negative design (TND) is a common and cost-efficient study design for estimating vaccine effectiveness [9, 10]. In its canonical form, patients with acute respiratory illness are prospectively screened and recruited based on a unified symptom set through hospitals or other care facilities and tested for the pathogen of interest. Vaccination history and other detailed covariate information are collected at enrollment or through linked medical or vaccination records. Effectiveness is then estimated by comparing the odds of vaccination among test-positive (cases) and test-negative (controls) individuals.
@@
The TND has been previously promoted as a way to reduce confounding due to unmeasured healthcare-seeking behavior — a common source of bias in observational evaluations of vaccines [10]. Because participants seek a test conditional on the same symptoms but without knowledge of the causative agent, the design effectively restricts to “care-seekers”, thereby precluding it as potential confounder [10]. Subsequent work pointed out that the TND (i) may not be robust to unmeasured healthcare-seeking when it is nonbinary or when there are other factors affecting vaccine uptake and risk of infection [11, 12, 13] and (ii) is subject to possible selection bias due to conditioning on receiving a test [11, 13].
More recently, [14] and [15] introduced a formal causal framework for the TND based on potential outcomes, allowing for a more precise discussion of identification of marginal and conditional vaccine effects in the TND. However, contrary to the original motivation for the TND, they require that observed covariates are sufficient to ensure no unmeasured confounding or selection bias. Thus, a question remains as to whether the TND may be formally justified under weaker assumptions.
@@
In this article, we revisit the original motivation and show that the TND may be justified under a less restrictive assumption that unmeasured confounding is equivalent for test-positive cases and test-negative controls on the odds ratio scale. This assumption builds on the insight that infection with another respiratory illness, if unaffected by focal vaccination, is essentially a negative outcome control [16, 17]. Therefore, any observed difference in the incidence of test negative illnesses between vaccinated and unvaccinated individuals likely reflects residual confounding. Under the assumption of confounding equivalence, we can use the difference in test-negative illness to de-bias our estimate of vaccine effectiveness, similar to the bias correction via the parallel trends assumption in the difference-in-differences literature [18, 19, 20]. As we show, this approach is valid under the outcome-dependent sampling of the TND, where only those who are tested are sampled. When our assumptions hold, we show that the traditional odds ratio estimator identifies the conditional causal risk ratio among the vaccinated. We also derive new estimators of the marginal risk ratio among the vaccinated, including those based on outcome modeling and inverse probability weighting. We discuss the implications of our assumptions for the design of TNDs and provide proofs of our results as well as simulations to illustrate the consequences when our assumptions are violated. We also propose a sensitivity analysis to evaluate the impact of deviations from equi-confounding on our effect estimates. Finally, although our focus is on the TND, we show how our bias adjustment approach under equi-confounding applies more broadly to cohort studies with symptom-triggered testing, i.e., where the outcome is only ascertained for those who seek a test after developing symptoms.
@@ 2 Setup and observed data
    the test is perfect such that $Y_{i}=2\implies I_{i}=2$ and $Y_{i}=1\implies I_{i}=1$ for every individual $i$.
These assumptions can be relaxed (e.g., we discuss imperfect testing in Appendix H). Under this testing regime, we have that $Y=I\times T$ and we consider $Y\neq 0$ to define medically-attended illness, i.e., symptomatic infection severe enough to warrant medical attention. Finally, let $Y^{v}$, $I^{v}$ and $T^{v}$ denote the potential outcomes of $Y$, II and $T$ respectively under $V=v$ for $v=0,1$.
@@ 3 Causal estimands for the test-negative design
In a TND study, the primary estimand is the marginal causal risk ratio,
@@
comparing the probability of medically-attended illness due to the pathogen of interest if everyone were vaccinated versus unvaccinated in the population underlying the TND, with causal vaccine effectiveness (VE) defined as 1RR1-$R$. Occasionally, the causal odds ratio is presented as the target parameter instead of the risk ratio. However, this is often combined with an assumption —either explicit or implicit— that symptomatic infection is rare, in which case the odds ratio approximates the risk ratio.
@@
may nevertheless be point identified. The parameter $\Psi$ is analogous to the average treatment effect on the treated (ATT) in the causal inference literature. Similarly, we can define a modified VE as $1-\Psi$, interpreted as the proportion of medically-attended illness prevented by vaccination among the vaccinated subjects. When protection is homogeneous, $\Psi$ coincides with the causal risk ratio, i.e. $RR=\Psi$; otherwise, when protection is heterogeneous, $RR$ is a weighted average of the causal risk ratios among the vaccinated and unvaccinated.
@@ 4 Illustrative example
To build intuition, we start with a brief account of the identification of the conditional risk ratio among the vaccinated, $\Psi(X)$, before turning to formal results for $\Psi$ in section 5.
@@ 4.1 Decomposition of the conditional causal risk ratio among the vaccinated
This decomposition expresses $\Psi(X)$ as the product of two terms: the (crude) observed risk ratio and a de-biasing term. The de-biasing term quantifies the degree of unmeasured confounding between the vaccinated and unvaccinated on the ratio scale, i.e., the net influence of all unmeasured causes of vaccination and the outcome. When there is no unmeasured confounding, the de-biasing term equals one and the observed risk ratio equals the causal risk ratio among the vaccinated. Conversely, in the presence of unmeasured confounding, the de-biasing term is the quantity necessary to remove bias from the observed risk ratio. However, without additional assumptions, this term (in particular the denominator) cannot be identified from the observed data.
@@ 4.3 Identification under cohort and outcome-dependent sampling
Nevertheless, under the invariance of the odds ratio [21], we can show that $\Psi_{om}(X)$ is equivalent to
@@
which is identified under the sampling design of the TND. As discussed further in section 6, $\Psi^{*}_{om}(X)$ is identical to the conditional odds ratio estimand commonly targeted in TND studies [10], suggesting that estimates from existing studies could be reinterpreted and justified under the equi-confounding conditions established here.
@@ 5.1 Identifiability conditions
    $\text{where }OR_{y}(X)=\frac{\Pr[Y^{0}=y|V=1,X]\Pr[Y^{0}=0|V=0,X]}{\Pr[Y^{0}=0|V=1,X]\Pr[Y^{0}=y|V=0,X]}.$
@@
Assumption A1 is a well-known identifiability condition discussed in more detail elsewhere [22]. Assumption A2 has two implications. First, it implies that the vaccine does not offer any cross-protection against other types of infection which may cause the same symptoms, including any short-term, nonspecific protection via activation of the immune system. Second, it requires that an individual’s decision to seek care and get tested is unaffected by their vaccination status, though it may depend on other measured and unmeasured factors. In theory, Assumption A2 could be evaluated in a randomized trial by comparing the incidence of testing negative (or testing positive for specific non-target pathogens) across arms. However, in practice, trials may not be adequately powered to rule out small effects. Both aspects of Assumption A2 have been invoked in previous literature on analysis of TND studies [10, 23, 14]. In Appendix G, we show that Assumption A2 can be further relaxed if the effect of vaccination on testing is equivalent for test-positive and test-negative illnesses and an alternative estimand is considered.
Assumption A3 is our key assumption and an alternative to the strict no unmeasured confounding (i.e. conditional independence) assumption suggested previously [14]. It states that the degree of unmeasured confounding on the odds ratio scale is the same for test-positive and test-negative illnesses. Notably, it does not require unmeasured confounders to be binary or even related to health-seeking specifically, although health-seeking behavior is likely a credible candidate for satisfying this condition. [24] mentioned a similar condition previously, although not in a formal causal framework [24]. The assumption is also similar to a recently suggested scale-independent alternative to the parallel trend assumption in the difference-in-differences literature [19, 20]. We discuss the underlying parameterization in more detail in Appendix A. In Appendix B, we also explore potential structural sources of equi-confounding —including health-seeking behavior and immortal time— as well as example violations, such as correlated vaccine-seeking [25] and at home testing [26] and how these may be resolved. When infections are mutually exclusive, Assumption A3 simplifies to risk ratio equivalence in Equation 2. Further, recalling $Y=I\times T$, we can split Assumption A3 into:
@@
Finally, Assumption A4 is similar to, although less restrictive than, the well-known positivity condition in causal inference [22]. Specifically, Assumption A4 requires overlap of vaccination propensity between the observed vaccinated and unvaccinated among the test positive cases as well as overlap of vaccination propensity across test-positive cases and test-negative controls within observed vaccination groups.
@@ 5.2 Identification in a cohort study
Proof of Proposition 1 is deferred to Appendix C.1. ∎
@@
Following standard terminology [27], we refer to Expression 6 as the outcome-modeling estimand and Expression 7 as the inverse-probability weighting estimand.
@@
The quantity $\Pr[V=1|Y^{0}=2,X]$ is sometimes referred to as the extended propensity score (EPS) function, as it extends the standard propensity score to allow for unmeasured confounding by conditioning on the treatment-free potential outcome — in this case $Y^{0}=2$. As shown previously [28], in the presence of unmeasured confounding, knowledge of the EPS is sufficient to identify treatment effects among the treated. Similarly knowledge of $\Pr[V=1|Y^{1}=2,X]$ would permit identification of treatment effects among the untreated, and knowledge of both would permit the identification of the overall effect [28].
@@
Recent work by [29] compares the performance of the TND with analyses in the full cohort using target trial emulation. However, it is unclear whether these designs target the same estimand. Indeed, Expressions 4, 6, and 7 suggest an alternative analysis in the cohort —essentially a difference-in-differences analysis comparing test-positive and test-negative outcomes— would target $\Psi$ and could offer a more direct comparison with the TND estimators introduced in this paper. We also note that, in a cohort study, alternative estimates of absolute risk differences could be derived under equi-confounding.
@@ 5.3 Identification under outcome-dependent sampling in a TND
Proof of Proposition 2 is deferred to Appendix C.2. ∎
@@
Expressions 8 and 9 resemble those proposed assuming unconfoundedness [15]. In that work, the authors introduce a class of marginal estimands in a TND that can be written as
@@ 6 Estimation
We now describe estimation strategies for the causal risk ratio among the vaccinated in a TND, based on the identifying expressions in Proposition 2. However, we describe in Appendix D estimators based on the results in Proposition 1 in a cohort study. For convenience, we define an indicator for test positive case in the TND sample $Y^{*}=\mathbbm{1}(Y=2)$.
@@
where $\pi_{y}(X)=\Pr[V=1\mid S=1,Y^{*}=y,X]$ is the probability of vaccination among the test-negative controls. This probability could be obtained, for instance, via a logistic regression of $V$ on $X$ among those with $Y^{*}=0$. In Appendix D, we discuss methods for estimating standard errors and confidence intervals for $\widehat{\Psi}_{om}^{*}$ and $\widehat{\Psi}_{ipw}^{*}$ via bootstrapping or stacked estimating equations.
The first estimator, $\widehat{\Psi}_{om}^{*}$, requires correct specification of the outcome model, while the second estimator, $\widehat{\Psi}_{ipw}^{*}$, requires correct specification of the extended propensity score model. In practice, however, both models may be difficult to specify correctly. To address this, we construct a doubly-robust estimator for $\Psi$ that is consistent and asymptotically normal if the model for either $\mu^{*}_{0}(X)$ or $\pi^{*}_{0}(X)$ or both are correctly specified. Full details of the doubly robust approach are included in Appendix F, but are based on noting that an equivalent expression of $\Psi_{om}$ and $\Psi_{ipw}$ is
@@
is the log of the conditional odds ratio function. As described in [30], a doubly-robust estimator of $\alpha^{*}(X)$ may be obtained as the solution to the empirical analogue of the population moment equation